# TokenRec: Learning to Tokenize ID for LLM-based Generative Recommendations

### v1

@@ 2.4.1. Tokenization&Prompts
2.4.2. User Modeling via LLM
The goal of user modeling component is to capture users’ preferences for generating the representations of items that user $u_{i}$ may like. A typical input $\mathcal{X}_{i}$ of our LLM backbone can be formed by selecting a prompt template $\mathcal{P}$ and the corresponding ID tokens for user $u_{i}$ and his/her interacted items $\mathcal{N}_{(u_{i})}$ in the history as follows:
(15) $\mathcal{X}_{i}\rightarrow(\mathcal{P},\mathcal{T}^{c}_{u_{i}})\ or\ (\mathcal{P},\mathcal{T}^{c}_{u_{i}},\{\mathcal{T}^{c}_{v_{j}}|v_{j}\in\mathcal{N}_{(u_{i})}\}),$
@@ 2.4.2. User Modeling via LLM
(16) $\mathcal{T}_{t}=\text{LLM}(\mathcal{X}_{i},\mathcal{T}_{i:t-1}),$
@@
(17) $\mathbf{h}_{i}=\text{LLM4Rec}(\mathcal{X}_{i}).$
@@
2.4.3. Generative Retrieval
In general, LLM-based recommender systems employ auto-regressive generation to decode recommendations in natural language (Geng et al., 2022; Zheng et al., 2023), such as producing textual strings like ”item_1234” or ”the user love electronics …”. However, the beam search decoding in LLMs can be very time-consuming during auto-regressive generation (Wang et al., 2024b), which is impractical for various real-time recommendation scenarios. Moreover, due to the hallucination issue, it is highly challenging to generate accurate item titles and descriptions when making personalized recommendations. For example, items’ title “iPhone SE, 256 GB, starlight” and “iPhone 15, 256 GB, starlight” share most tokens but are significantly different products - with “iPhone 15, 256 GB, starlight’ being a non-factual product. The hallucination issue is likely more severe in e-commerce platforms, in which billions of products are sold, leading to invalid item identifiers for recommendations. Furthermore, title generation in LLM-based recommendations cannot generate unseen items in the fine-tuning stage, which is often infeasible in practice. To this end, we propose a generative retrieval paradigm for LLM-based recommendations, where a simple but effective and efficient strategy is designed to project generative users’ preferences for retrieving potential items from the whole item pool.
@@ 2.4.3. Generative Retrieval
(18) $\mathbf{z}_{i}=\text{Proj}(\mathbf{h}_{i}),$
where $\text{Proj}(\cdot)$ can be modeled by a three-layer MLPs. Note that $\mathbf{z}_{i}$ can be considered the generative representation of the next recommended items for user $u_{i}$’. After that, we propose to retrieve the $K$-nearest items from the whole item set $\mathcal{V}$ for generating the personalized top-$K$ recommendation list. This can be achieved by measuring the similarity scores between the target user’s generative preference and high-order items’ representations from the well-trained GNNs-based recommendation method. For example, the predicted similarity score $y_{ij}$ of user $u_{i}$ towards item $v_{j}$ can be calculated by a matching function (e.g., cosine similarity) between user $u_{i}$’s generative item representation $\mathbf{z}_{i}$ and GNN-based item’s $v_{j}$ representation $\mathbf{q}_{j}$ which is stored in a vector database:
(19) $y_{ij}=\frac{\mathbf{z}_{i}\mathbf{q}_{j}}{\|\mathbf{z}_{i}\|\|\mathbf{q}_{j}\|}.$
@@
2.5. TokenRec’s Training and Inference
2.5.1. Training
@@
  • Step 2. Tuning the LLM4Rec for Generative Retrieval. In this step, we tune the LLM backbone (e.g., T5), LLM token embeddings, and the projection layer for generative retrieval recommendations while keeping the MQ-Tokenizers frozen. More specifically, the objective of generative retrieval is to identify potential items from the whole item pool, such that top-$K$ ranked items are relevant to the specific query (i.e., generative item representations of users). To achieve such an objective, a general solution is to perform metric learning by predicting the relative similarity or distance based on the dense representations between inputs (Cakir et al., 2019; McFee and Lanckriet, 2010; Liu et al., 2009). In other words, the proposed TokenRec aims to perform the nearest neighbor retrieval, achieved by calculating ranking scores between the user’s generative item representation and collaborative item representations learned from GNNs, so as to retrieve top-$K$ next items to the target user for personalized recommendations. Mathematically, we utilize a pairwise ranking loss for our tuning process, which encourages the query closer to the positive set (i.e., when $\lambda=1$) than to the negative set by a fixed margin $\gamma$ (i.e., when $\lambda=-1$) as follows:
    (20) $\mathcal{L}_{\text{LLM4Rec}}=\begin{cases}1-\text{sim}(\mathbf{z}_{i},\mathbf{q}_{j}),&\text{if}\ \lambda=1\\ \max(0,\text{sim}(\mathbf{z}_{i},\mathbf{q}_{j})-\gamma),&\text{if}\ \lambda=-1\end{cases}$
@@
Figure 3. The TokenRec’s efficiency and generalization capability for new users and items during the inference stage. Rather than retraining the entire model, which can be computationally expensive and time-consuming, only the GNN needs to be updated for learning representations for new users and items while keeping the MQ-Tokenizers and LLM backbone frozen. This significantly reduces the computational burden and allows faster adaptation to new users and items.
2.5.2. Inference
@@
  • Efficient Recommendations. TokenRec proposes a novel LLM-empowered collaborative recommendation framework in generative retrieval paradigms, bypassing the time-consuming decoding process. In particular, the proposed LLM-based recommendation framework aims to output generative item representation and retrieve appropriate items for collaborative recommendations instead of a sequence of discrete tokens (e.g., item titles). This contributes to reducing the considerable computing cost of online RecSys. The efficiency evaluation can be found in Section 3.4 and Table 5.
@@
  • Generalizability to New Users and Items. By employing masking and $K$-way encoder mechanisms in vector quantization, the proposed architecture can provide robust ID tokenization for unseen users and items in fine-tuning the LLM4Rec component. For instance, as shown in Figure 3, when new users and items are added to the system, retraining is only required for the collaborative filtering component (e.g., MF and LightGCN) to capture collaborative knowledge in learning representations of new users and items (i.e., updating vector database). The MQ-Tokenizers and LLM backbone can remain frozen, saving computational resources and eliminating the need for retraining these components. Compared to finetuning LLMs, the training of GNNs is far more efficient (Zhang et al., 2024). The capability in adapting to new users and items is demonstrated in Section 3.3 and Table 4.
@@
  • Concise Prompts. TokenRec provides an inference alternative that relies solely on user ID tokens, e.g., Prompt 1 in Section 2.4.1, for LLM-based recommendation generation. This is made possible by incorporating the collaborative knowledge of users into user ID tokens through our MQ-Tokenizer. By doing so, TokenRec eliminates the necessity of including interacted items in inputs, thus reducing significant computing resources during inference. Additionally, this mechanism proves advantageous when dealing with users who have interacted with a large number of items, effectively avoiding the prevalent issue of context length limitation of many LLMs (Liu et al., 2024; Kaneko and Okazaki, 2023), e.g., 512 tokens in T5 (Raffel et al., 2020) and 2048 tokens in ChatGPT. The recommendation performance using user ID tokens only for model input can be observed in Table 2 and Table 3.
3. Experiment
3.1. Experimental Settings
3.1.1. Datasets
To demonstrate the effectiveness of the proposed method, we conduct comprehensive experiments over four benchmark datasets: Amazon-Beauty (Beauty for short), Amazon-Clothing (Clothing for short), LastFM, and MovieLens 1M (ML1M for short). The first two datasets [^2] https://nijianmo.github.io/amazon/ are obtained from the amazon.com e-commerce platform, encompassing a broad spectrum of user interactions with Beauty and Clothing products. The LastFM [^3] https://grouplens.org/datasets/hetrec-2011/ dataset provides music artist listening records from users at the Last.fm online music system. The ML1M [^4] https://grouplens.org/datasets/movielens/1m/ dataset offers a collection of movie ratings made by MovieLens users. Table 1 shows the statistics of these four datasets, where the maximum item sequence length is set to 100 to accommodate the input length of the LLM backbone T5 (512 tokens). Moreover, our training-validation-testing combination follows the leave-one-out policy (Geng et al., 2022), i.e., using all but the last observation in users’ interaction history as the training set. Last but not least, we randomly shuffle users’ interaction history to provide a non-chronological item list to align with the settings of collaborative filtering methods.
Table 1. Basic statistics of benchmark datasets.
@@ 3.1.1. Datasets
3.1.2. Baselines
@@
  • Collaborative Filtering: MF (Rendle et al., 2009) is the most classic CF method, while NeuCF (He et al., 2017) is the very first DNNs based collaborative filtering. LightGCN (He et al., 2020) and GTN (Fan et al., 2022) are representative CF based on GNNs techniques. LTGNN (Zhang et al., 2024) is the most advanced GNNs-based collaborative filtering method.
@@
  • Sequential Recommendations: SASRec (Kang and McAuley, 2018) is a self-attention based sequential recommendation model. BERT4Rec (Sun et al., 2019) is a bidirectional Transformer-based recommender trained with the BERT-style cloze task. S3Rec (Zhou et al., 2020) is a representative sequential recommendation model trained by self-supervised learning.
@@
  • P5 (Geng et al., 2022) is a pioneering work on LLM-based RecSys, which describes recommendation tasks in a text-to-text format and employs LLMs to capture deeper semantics for personalization and recommendation. In our experiments, we deploy two indexing methods, i.e., random indexing (RID) and sequential indexing (SID), on the P5 model. Among these, P5-SID is chosen as a baseline for our efficiency evaluation, generalizability evaluation, and ablation study in our experiments.
@@
  • CID (Hua et al., 2023) is a non-trivial indexing approach that considers the co-occurrence matrix of items to design numeric IDs so that items co-occur in user-item interactions will have similar numeric IDs. To be consistent, we employ the P5 model as its LLM backbone.
@@
  • POD (Li et al., 2023) encodes discrete prompts into continuous embeddings to reduce the excessive input length of LLMs based on P5 architecture.
@@
  • CoLLM (Zhang et al., 2023) employs GNNs to provide continuous embeddings representing items and users for LLM-based recommendations. Notably, in our experiments, CoLLM’s binary classification output (i.e., whether or not a user likes an item) is reformatted to generate item IDs consistent with the output setting of top-$K$ recommendations.
3.1.3. Evaluation Metrics
@@
In order to evaluate the quality of recommendation results, we adopt two widely used evaluation metrics: top-$K$ Hit Ratio (HR@$K$) and top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$(Wang et al., 2019; Zhang et al., 2024), in which higher values indicate better performance for recommendations. And the average metrics for all users in the test set are reported. In addition, we set the values of $K$ as 10, 20, and 30, among which 20 is the default value for ablation experiments.
3.1.4. Hyper-parameter Settings
@@
Our proposed model is implemented based on Hugging Face and PyTorch. The codebook number $K$ (i.e., the subencoder), the token number $L$ at each subcodebook, and the ratio $\rho$ of our masking operation are search in the ranges of $\{1,2,3,4,5\}$, $\{128,256,512,1024\}$, and {0 to 1} in 0.1 increments, respectively. Moreover, during the fine-tuning process of our proposal LLM-based recommendation framework, the ratio of negative sampling $\lambda$ presented in Eq.(20) is fixed at 1:1, in which we randomly select an un-interacted item from the whole item base as the negative sample for each positive sample. The margin $\gamma$ in Eq.(20) is set to 0 to 0.2. We optimize the MQ-Tokenizers and the LLM backbone with AdamW (Loshchilov and Hutter, 2019) in a mini-batch manner. Note that the high-order collaborative representations for users and items are obtained from a representative collaborative filtering method, namely LightGCN (He et al., 2020). For prompting, we design 11 templates for TokenRec: 10 of them are set to be seen prompts, and the remaining one is evaluated as an unseen prompt. For a fair comparison, we employ a widely-used lightweight LLM, i.e., T5-small (Raffel et al., 2020), for our LLM4Rec model and all LLM-based baselines. The other default hyper-parameters for baseline methods are set as suggested by the corresponding papers.
Table 2. Performance comparison of recommendation algorithms on the LastFM and ML1M datasets.
@@
Table 3. Performance comparison of recommendation algorithms on the Beauty and Clothing datasets.
@@
3.2. Performance Comparison of Recommender Systems
We first compare the recommendation performance between TokenRec and all baselines over four benchmark datasets. Table 2 and Table 3 present the overall performance comparison on the four datasets, where * denotes the proposed TokenRec, and the best and second best results are marked by Bold and underlined. Notably, the proposed model using unseen prompts will be indicated by the suffix (Unseen Prompt). Besides, the suffix of (User ID Only) denotes the case of using user ID tokens only for LLM-empowered recommendation without considering users’ interaction history towards items. We make the following observations:
@@ 3.2. Performance Comparison of Recommender Systems
  • As an earlier non-trivial indexing method to capture hidden knowledge in co-occurrence frequency, Collaborative Indexing (CID) outperforms Random Indexing (RID) and Sequential Indexing (SID) given the same setting of P5. These observations suggest the potential of integrating collaborative knowledge for item&user tokenization/indexing. However, the P5 variants and POD are inferior to the existing advanced GNN-based collaborative filtering (i.e., LightGCN, GTN, and LTGNN), implying their inability to capture collaborative information using LLMs.

### v2

@@ II-D1 Tokenization&Prompts
II-D2 User Modeling via LLM
The goal of the user modeling component is to capture users’ preferences for generating the representations of items that user $u_{i}$ may like. A typical input $\mathcal{X}_{i}$ of our LLM backbone can be formed by selecting a prompt template $\mathcal{P}$ and the corresponding ID tokens for user $u_{i}$ and his/her interacted items $\mathcal{N}_{(u_{i})}$ in the history as follows:
$\mathcal{X}_{i}\rightarrow(\mathcal{P},\mathcal{T}^{c}_{u_{i}})\ or\ (\mathcal{P},\mathcal{T}^{c}_{u_{i}},\{\mathcal{T}^{c}_{v_{j}}|v_{j}\in\mathcal{N}_{(u_{i})}\}),$ (15)
@@ II-D2 User Modeling via LLM
$\mathcal{T}_{t}=\text{LLM}(\mathcal{X}_{i},\mathcal{T}_{i:t-1}),$ (16)
@@
$\mathbf{h}_{i}=\text{LLM4Rec}(\mathcal{X}_{i}).$ (17)
@@
II-D3 Generative Retrieval
In general, LLM-based recommender systems employ auto-regressive generation to decode recommendations in natural language [10, 24], such as producing textual strings like ”item_1234” or ”the user love electronics …”. However, the beam search decoding in LLMs can be very time-consuming during auto-regressive generation [25], which is impractical for various real-time recommendation scenarios. Moreover, due to the hallucination issue, generating accurate item titles and descriptions is highly challenging when making personalized recommendations. For example, items’ title “iPhone SE, 256 GB, starlight” and “iPhone 15, 256 GB, starlight” share most tokens but are significantly different products - with “iPhone 15, 256 GB, starlight’ being a non-factual product. The hallucination issue is likely more severe in e-commerce platforms, where billions of products are sold, leading to invalid item identifiers for recommendations. Furthermore, title generation in LLM-based recommendations cannot generate unseen items in the fine-tuning stage, which is often infeasible in practice. To this end, we propose a generative retrieval paradigm for LLM-based recommendations, where a simple but effective and efficient strategy is designed to project generative users’ preferences for retrieving potential items from the whole item pool.
@@ II-D3 Generative Retrieval
$\mathbf{z}_{i}=\text{Proj}(\mathbf{h}_{i}),$ (18)
where $\text{Proj}(\cdot)$ can be modeled by a three-layer MLPs. Note that $\mathbf{z}_{i}$ can be considered the generative representation of the next recommended items for user $u_{i}$’. After that, we propose to retrieve the $K$-nearest items from the whole item set $\mathcal{V}$ for generating the personalized top-$K$ recommendation list. This can be achieved by measuring the similarity scores between the target user’s generative preference and high-order items’ representations from the well-trained GNN-based recommendation method. For example, the predicted similarity score $y_{ij}$ of user $u_{i}$ towards item $v_{j}$ can be calculated by a matching function (e.g., cosine similarity) between user $u_{i}$’s generative item representation $\mathbf{z}_{i}$ and GNN-based item’s $v_{j}$ representation $\mathbf{q}_{j}$ which is stored in a vector database:
$y_{ij}=\frac{\mathbf{z}_{i}\mathbf{q}_{j}}{\|\mathbf{z}_{i}\|\|\mathbf{q}_{j}\|}.$ (19)
@@
II-E TokenRec’s Training and Inference
II-E1 Training
@@
  • Step 2. Tuning the LLM4Rec for Generative Retrieval. In this step, we tune the LLM backbone (e.g., T5), LLM token embeddings, and the projection layer for generative retrieval recommendations while keeping the MQ-Tokenizers frozen. More specifically, the objective of generative retrieval is to identify potential items from the whole item pool, such that top-$K$ ranked items are relevant to the specific query (i.e., generative item representations of users). To achieve such an objective, a general solution is to perform metric learning by predicting the relative similarity or distance based on the dense representations between inputs [26, 27, 28]. In other words, the proposed TokenRec aims to perform the nearest neighbor retrieval, achieved by calculating ranking scores between the user’s generative item representation and collaborative item representations learned from GNNs, so as to retrieve top-$K$ next items to the target user for personalized recommendations. Mathematically, we utilize a pairwise ranking loss for our tuning process, which encourages the query closer to the positive set (i.e., when $\lambda=1$) than to the negative set by a fixed margin $\gamma$ (i.e., when $\lambda=-1$) as follows:
    $\mathcal{L}_{\text{LLM4Rec}}=\begin{cases}1-\text{sim}(\mathbf{z}_{i},\mathbf{q}_{j}),&\text{if}\ \lambda=1\\ \max(0,\text{sim}(\mathbf{z}_{i},\mathbf{q}_{j})-\gamma),&\text{if}\ \lambda=-1\end{cases}$ (20)
@@
Figure 3: The TokenRec’s efficiency and generalization capability for new users and items during the inference stage. Rather than retraining the entire model, which can be computationally expensive and time-consuming, only the GNN needs to be updated for learning representations for new users and items while keeping the MQ-Tokenizers and LLM backbone frozen. This significantly reduces the computational burden and allows faster adaptation to new users and items.
II-E2 Inference
@@
  • Efficient Recommendations. TokenRec proposes a novel LLM-empowered collaborative recommendation framework in generative retrieval paradigms, bypassing the time-consuming decoding process. In particular, the proposed LLM-based recommendation framework aims to output generative item representation and retrieve appropriate items for collaborative recommendations instead of a sequence of discrete tokens (e.g., item titles). This contributes to reducing the considerable computing cost of online RecSys. The efficiency evaluation can be found in Section III-D and Table V.
@@
  • Generalizability to New Users and Items. The proposed architecture can provide robust ID tokenization for unseen users and items without fine-tuning the LLM4Rec component. As shown in Figure 3, when new users and items are added to the system, retraining is only required for the collaborative filtering component (e.g., MF and LightGCN) to capture collaborative knowledge in learning representations of these users and items (i.e., updating vector database). In contrast, the MQ-Tokenizers and LLM backbone can remain frozen and perform well, thanks to the masking and $K$-way encoder mechanisms in vector quantization. This eliminates the need for retraining LLM-related components and therefore saves huge computational resources: compared to the finetuning of LLMs, the training of GNNs is far more efficient [29]. Notably, the capability to adapt to new users and items is demonstrated in Section III-C and Table IV.
@@
  • Concise Prompts. TokenRec provides an inference alternative that relies solely on user ID tokens, e.g., Prompt 1 in Section II-D1, for LLM-based recommendation generation. This is made possible by incorporating the collaborative knowledge of users into user ID tokens through our MQ-Tokenizer. By doing so, TokenRec eliminates the necessity of including interacted items in inputs, thus reducing significant computing resources during inference. Additionally, this mechanism proves advantageous when dealing with users who have interacted with a large number of items, effectively avoiding the prevalent issue of context length limitation of many LLMs [30, 31], e.g., 512 tokens in T5 [32] and 2048 tokens in ChatGPT. The recommendation performance using user ID tokens only for model input can be observed in Table II and Table III.
III Experiment
III-A Experimental Settings
III-A1 Datasets
[…]
TABLE I: Basic statistics of benchmark datasets.
@@ III-A1 Datasets
  • # represents the number of users, items, and interactions.
III-A2 Baselines
@@
  • Collaborative Filtering: MF [33] is the most classic CF method, while NeuCF [34] is the very first DNNs based collaborative filtering. LightGCN [5] and GTN [6] are representative CF based on GNNs techniques. LTGNN [29] is the most advanced GNN-based collaborative filtering method.
@@
  • Sequential Recommendations: SASRec [35] is a self-attention based sequential recommendation model. BERT4Rec [36] is a bidirectional Transformer-based recommender trained with the BERT-style cloze task. S3Rec [37] is a representative sequential recommendation model trained by self-supervised learning.
@@
  • P5 [10] is a pioneering work on LLM-based RecSys, which describes recommendation tasks in a text-to-text format and employs LLMs to capture deeper semantics for personalization and recommendation. In our experiments, we deploy two indexing methods, i.e., random indexing (RID) and sequential indexing (SID), on the P5 model. Among these, P5-SID is chosen as a baseline for our efficiency evaluation, generalizability evaluation, and ablation study in our experiments.
@@
  • CID [38] is a non-trivial indexing approach that considers the co-occurrence matrix of items to design numeric IDs so that items co-occur in user-item interactions will have similar numeric IDs. To be consistent, we employ the P5 model as its LLM backbone.
@@
  • POD [39] encodes discrete prompts into continuous embeddings to reduce the excessive input length of LLMs based on P5 architecture.
@@
  • CoLLM [14] employs GNNs to provide continuous embeddings representing items and users for LLM-based recommendations. Notably, in our experiments, CoLLM’s binary classification output (i.e., whether or not a user likes an item) is reformatted to generate item IDs consistent with the output setting of top-$K$ recommendations.
III-A3 Evaluation Metrics
@@
In order to evaluate the quality of recommendation results, we adopt two widely used evaluation metrics: top-$K$ Hit Ratio (HR@$K$) and top-$K$ Normalized Discounted Cumulative Gain (NDCG@$K$[18, 29], in which higher values indicate better performance for recommendations. And the average metrics for all users in the test set are reported. In addition, we set the values of $K$ as 10, 20, and 30, among which 20 is the default value for ablation experiments.
III-A4 Hyper-parameter Settings
@@
Our proposed model is implemented based on Hugging Face and PyTorch. The codebook number $K$ (i.e., the subencoder), the token number $L$ at each subcodebook, and the ratio $\rho$ of our masking operation are search in the ranges of $\{1,2,3,4,5\}$, $\{128,256,512,1024\}$, and {0 to 1} in 0.1 increments, respectively. Moreover, during the fine-tuning process of our proposal LLM-based recommendation framework, the ratio of negative sampling $\lambda$ presented in Eq.(20) is fixed at 1:1, in which we randomly select an un-interacted item from the whole item base as the negative sample for each positive sample. The margin $\gamma$ in Eq.(20) is set to 0 to 0.2. We optimize the MQ-Tokenizers and the LLM backbone with AdamW [40] in a mini-batch manner. Note that the high-order collaborative representations for users and items are obtained from a representative collaborative filtering method, namely LightGCN [5]. For prompting, we design 11 templates for TokenRec: 10 of them are set to be seen prompts, and the remaining one is evaluated as an unseen prompt. For a fair comparison, we employ a widely-used lightweight LLM, i.e., T5-small [32], for TokenRec and all LLM-based baselines. The other default hyper-parameters for baseline methods are set as suggested by the corresponding papers.
TABLE II: Performance comparison of recommendation algorithms on the LastFM and ML1M datasets.
@@
TABLE III: Performance comparison of recommendation algorithms on the Beauty and Clothing datasets.
@@
III-B Performance Comparison of Recommender Systems
We first compare the recommendation performance between TokenRec and all baselines over four benchmark datasets. Table II and Table III present the overall performance comparison on the four datasets, where * denotes the proposed TokenRec, and the best and second best results are marked by Bold and underlined. Notably, the proposed model using unseen prompts will be indicated by the suffix (Unseen Prompt). Besides, the suffix of (User ID Only) denotes the case of using user ID tokens only for LLM-empowered recommendation without considering users’ interaction history towards items. We make the following observations:
@@ III-B Performance Comparison of Recommender Systems
  • As an earlier non-trivial indexing method to capture hidden knowledge in co-occurrence frequency, Collaborative Indexing (CID) outperforms Random Indexing (RID) and Sequential Indexing (SID) given the same setting of P5. These observations suggest the potential of integrating collaborative knowledge for item&user tokenization/indexing. However, the P5 variants and POD are inferior to the existing GNN-based collaborative filtering (i.e., LightGCN, GTN, and LTGNN), implying their inability to capture collaborative information using LLMs.